OpenStudy (anonymous):

need help finding the gen. sol. of the differential equation (D-1)^2 (D+3)(D^2 + 2D + 5)^2[y] = 0

OpenStudy (lgbasallote):

this is \[\huge [(D-1)^2 (d+3)(D^2 + 2D +5)^2] y = 0\] right?

OpenStudy (lgbasallote):

interesting

OpenStudy (lgbasallote):

so what would be the problem?

OpenStudy (anonymous):

how would i turn those back into y-primes?

OpenStudy (lgbasallote):

why would you want to turn them into y primes?

OpenStudy (anonymous):

good question. i shouldnt

OpenStudy (anonymous):

in the form of x in a general solution

OpenStudy (lgbasallote):

have you forgotten that D = m? you can just substitute straight to m \[\huge [(m-1)^2 ( m+3)(m^2 + 2m + 5)^2]y = 0\]

OpenStudy (lgbasallote):

can you see it now?

OpenStudy (anonymous):

i see. I will try to get it. I just dont like the looks of two of them being squared

OpenStudy (lgbasallote):

haha it's no worries...it just means the root has multiplicity 2

OpenStudy (anonymous):

would you mind showing me how it is done for the (m-1)^2? That will give me a great base for the long square

OpenStudy (lgbasallote):

i'll just give an example suppose you have \[\huge (m+3)^2 (m-4) = 0\] the roots are m = -3 ( 2x) and m=4 so the general solution is \[\huge y = y_c = c_1 e^{-3x} + c_2 x e^{-3x} + c_3 e^{4x}\]

OpenStudy (lgbasallote):

did that help?

OpenStudy (anonymous):

daaaamn this is a long problem

OpenStudy (anonymous):

yes. thank you again

OpenStudy (lgbasallote):

hahaha welcome ^_^

OpenStudy (lgbasallote):

and it's not really a long problem...just do q.f. in that long trinomial

OpenStudy (anonymous):

more complex solutions yay...

OpenStudy (lgbasallote):

haha